OpenStudy (anonymous):

Write a stigma notation for this series: (1/3)+(1/6)+(1/9)+(1/12)+....

OpenStudy (anonymous):



OpenStudy (sidsiddhartha):

the series is reciprocal of 3's multiples is'n it?

OpenStudy (sidsiddhartha):

so u can write it like\[\frac{ 1 }{ 3 }+\frac{ 1 }{ 3 \times 2 }+\frac{ 1 }{ 3 \times 3 }+\frac{ 1 }{ 3 \times 4 }+.......+\frac{ 1 }{ 3 \times ~n }\]

OpenStudy (sidsiddhartha):

ok?

OpenStudy (anonymous):

oh ok thanks for your help

OpenStudy (sidsiddhartha):

finally u can write it as\[\Large \sum_{n=1}^{\infty}\frac{ 1 }{ 3n }\]

OpenStudy (sidsiddhartha):

:)